# Enumerative Geometry of Quantum Periods

### v1

@@ 2.2. Definition of the mirror curve
We view the mirror curve as being fibered over Spec [$\text{Spec }\mathbb{C}[\textup{NE}(\mathbb{P})]$ via the closed moduli $z^{\beta}$, and we write $p(x,y,z)$. Write $p(x,y,z)=1-f(x,y,z)$. We will give an explicit definition of the mirror curve by specifying $f$.
@@
$\sum_{i=1}^{r+2}Q_{ik}m_{k}=0,\quad j=1,\ldots,r,$
can be packed into the charge matrix $Q=(Q_{ik})_{i,k}$. Using Gauß elimination we can transform the first $r$ columns of $Q$ into the identity matrix, so that $Q_{ik}=\delta_{ik}$ for $i,k\leq r$. Since $\mathbb{P}_{\Sigma}$ is smooth, the ray generators of $\rho_{r+1}$ and $\rho_{r+2}$ span the lattice, so all entries of $Q$ are integers. The rows of $Q$ define generators $\beta_{1},\ldots,\beta_{r}$ of the group $H_{2}(\mathbb{P}_{\Sigma},\mathbb{Z})$ by intersections with toric divisors
@@
Here we consider the indices modulo $r+2$. For an integer $l\geq 0$ and ordered tuples of integers $0\leq i_{1}<\ldots<i_{l}\leq m$ and $0\leq j_{1}<\ldots<j_{l}\leq n$ define
$\beta_{k,(i_{1},\ldots,i_{l}),(j_{1},\ldots,j_{l})}=\beta_{k}+\sum_{p=1}^{l}\sum_{q=0}^{i_{p}-1}\tilde{\beta}_{k-q}+\sum_{p=1}^{l}\sum_{q=0}^{j_{p}-1}\tilde{\beta}_{k+q}.$
@@
Define the toric potential of $\mathbb{P}_{\Sigma}$, resp. $\mathbb{P}$, by
@@
    The relation obtained from specializing the closed moduli $z^{\beta}$ by pairing with the polarization was already known by [CPS], Theorem 8.2, which amounts to saying that $\binom{\ell}{m_{k}}=\binom{\ell}{n_{k}}$ counts the summands in the $x^{a_{k}}y^{b_{k}}$-coefficient of $\vartheta_{1}$. Lemma 3.6 extends the agreement in part (1) of the remark by giving the correct closed moduli.
@@
Consider $\mathbb{P}^{1}\times\mathbb{P}^{1}$. A fan $\Sigma$ for $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is generated by $(-1,0)$, $(0,-1)$, $(1,0)$ and $(0,1)$. Again, all generators are vertices of $\Delta^{\star}$. The construction above gives
@@
Again consider $\mathbb{P}^{1}\times\mathbb{P}^{1}$, but this time with a $\mathbb{Q}$-Gorenstein degeneration to the Hirzebruch surface $\mathbb{P}_{\Sigma}=\mathbb{F}_{2}$. The fan $\Sigma$ is generated by $(-1,0)$, $(-2,-1)$, $(1,0)$ and $(0,1)$. The construction above gives
@@ 2.4. The quantum A-period
According to [CKYZ], §6, one (classical) period to consider is the integral of $\log(p)$ over the torus $|x|=|y|=1.$ We call this the A-period. Note that with $p=-f(x,y,z)$ we have $\log(p)=-\sum_{n\geq 1}\tfrac{1}{n}f^{n}$, and each term $f^{n}$ is a Laurent polynomial in $x$ and $y$. Recall from §1.2 that for such Laurent series we write ${\sf Const}(F)$ for the constant term — then the integral picks up the constant term ${\sf Const}(\log(p))=-\sum_{n\geq 1}\tfrac{1}{n}{{\sf Const}}(f^{n})$. We use this fact to generalize the definition of period to the non-commutative case, with non-commuting variables $X$ and $Y$ obeying $YX=qXY.$
@@ 2.5. Wavefunctions and the A-period
It is shown in [ACDKV], Section 2.2 that the one-form $\log Yd\log X$ acquires a pole at the point of brane insertion with residue $\hbar$ (recall $\hbar=g_{s}=\lambda$ is the string coupling) – equivalently $W$ changes by $\hbar$ times a period.222 Quantum mirror curves play a central role in the Topological Strings/Spectral Theory duality of [GHM], where the spectrum of the operator defining the mirror curve is studied. Further period computations can be found in many other works — see, e.g., [DH, HMMO, HK, HW, DL], for example.
@@
See §2.6 for the computation of $Y(X)$ for $(\mathbb{P}^{2},E)$. We make no claims about the existence of other solutions.
@@
Let $Y(X)$ be a solution to $p(X,Y)=0$ and let $\Psi(X)$ be the wavefunction defined by $p\cdot\Psi=0$. Then
@@ 2.7. Equality of definitions for the quantum A-period
Recall from §2.2 that $p(X,Y)=1-f(X,Y)$ gets quantized to an operator which acts on a function $\Phi(X)$ by $X\cdot\Phi(X)=X\Phi(X)$ and $Y\cdot\Phi(X)=\Phi(qX)$, and the wavefunction $\Psi(X)$ is defined by $p\cdot\Psi=0$. We have defined
@@
Substituting $X=xt$ and $Y=yt$, we get $f(X,Y,z)=tf(x,y,z)|_{z^{\beta}=t^{\beta\cdot E}z^{\beta}}$. E.g., for $\mathbb{P}^{2}$ we have $f=X+Y+zX^{-1}Y^{-1}=t(x+y+t^{3}zx^{-1}y^{-1})$. Applying $D$ as in §1.2, the claim is equivalent to
@@
Note that $g(t,y):=(1-tf)^{-1}\in(R[y]\llbracket y^{-1}\rrbracket)\llbracket t\rrbracket$ for $R=\mathbb{C}[\textup{NE}(X)][x]\llbracket x^{-1}\rrbracket$. By Convention 2.5, we can assume that $f(X,Y)$ does not contain any terms of $Y$-order $>1$. In other words, $\operatorname{ord}_{Y}(f)=1$, with Definition A.2. By Lemma A.3, $y(x,f)$ is a Laurent series of order $1$ in $f$. Then we can use Lemma A.6 to get
@@ 2.8. Natural $q$-refinements of mirror maps
where the last equality holds for every non-zero class $\beta\in\operatorname{NE}(\mathbb{P})$. The series $M(Q,q)$ will be given an enumerative interpretation in Theorem 4.2.
@@ 3. Quantum periods, broken lines and descendant invariants
In this section, we give a combinatorial construction of the quantum period (Definition 2.14) via broken lines in a consistent scattering diagram. In §4 we will use this description to prove that the primitive theta function, defined using broken lines, equals the open mirror map, also after $q$-refinement. Moreover, we discuss a relationship with descendant Gromov-Witten invariants with $\lambda$-class insertions.
@@ 3.1. Toric degenerations and scattering diagrams
Let $\mathbb{P}$ be a smooth projective Fano surface and let ${E}$ be a smooth anticanonical divisor, so that $(\mathbb{P},{E})$ is a smooth log Calabi-Yau pair. As in §2.1, if $\mathbb{P}$ is not toric, consider a $\mathbb{Q}$-Gorenstein degeneration of $\mathbb{P}$ to a Gorenstein toric Fano variety $\mathbb{P}_{\Sigma^{\prime}}$ or a smooth toric semi-Fano variety $\mathbb{P}_{\Sigma}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the ray generators of $\Sigma^{\prime}$ (or $\Sigma$) and let $\Delta$ be the polar dual polytope. Then $\Delta$ is the toric polytope of $\mathbb{P}_{\Sigma^{\prime}}$. The lattice points of $\Delta$ define an embedding of the projective toric variety $\mathbb{P}_{\Sigma^{\prime}}$ into $\mathbb{P}^{|\Delta\cap\mathbb{Z}^{2}|-1}$. The polytope $\Delta$ is reflexive. In particular it contains a unique interior lattice point.
@@ 3.2. Theta functions and quantum periods
Recall from §1.2 that we defined an operation $D$ on $\mathbb{C}[\textup{NE}(\mathbb{P})]$ by $Dz^{\beta}=(\beta\cdot D)z^{\beta}$.
@@
where $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ is the tropical analogue of $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$, defined as a count (with $q$-refined multiplicity) of rational tropical curves on the dual intersection complex $B$ of $(\mathbb{P},E)$ with an unbounded leg of weight $\beta\cdot E-1$ through a prescribed point, an unbounded leg of weight $1$ and no other unbounded legs. By [Gra2], Theorem 6.2, we have
@@
For $(\mathbb{P}^{2},E)$, Figure 3.3 shows the broken lines that give
@@ 3.3. Quantum descendant invariants
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the Chern classes of the Hodge bundle.
@@ 4. The $q$-refined open mirror map is the quantum theta function
We prove the identification of the quantum theta function asymptotic broken line expansion and the q-refined open mirror map in this section. As a consequence, we obtain the enumerative expression for the quantum period.
@@
Let $A=\bigoplus_{k\geq 0}A_{k}$ be a finitely generated graded algebra over $A_{0}$ and let $\widehat{A}$ denote the formal completion of $A$ in the ideal $\bigoplus_{k\geq 1}A_{k}$. Let $\theta(y)\in A\llbracket y^{-1}\rrbracket[y]$ be a formal Laurent series of the form $\theta(y)=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$. The series $\widetilde{y}=-\exp\left({-\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]\theta^{k}}\right)$ gives a well-defined element of $\widehat{A}$ that satisfies
$\theta(\tilde{y})=-1.$
Moreover, $\tilde{y}$ uniquely determines the coefficients $a_{k}$, hence it determines a function $\theta(y)$ of the given form.
@@
Define $f(y)=\theta(y^{-1})$ and let $A_{r}$ be the degree $r$ graded piece of $A$. Then $f(y)$ is a formal Laurent series in $y$ of the form $f(y)=y^{-1}+\sum_{k\geq 0}f_{k}y^{k}$ such that $f_{k}\in A_{k+1}$. Recall from §1.2 that we write $[y^{k}]f$ for the coefficient of $y^{k}$ in $f$. The power series $h(y):=yf(y)=1+\sum_{k>0}f_{k-1}y^{k}\in A\llbracket y\rrbracket$ has the convenient property $[y^{k}]h^{l}\in A_{k}$ for every $l\geq 0$. In particular, $[y^{k}](yf)^{k}\in A_{k}$, so $[y^{0}]f^{k}\in A_{k}$ and
@@
We apply Proposition A.12 to $h$ to have
@@
Comparing the very last expression with Lemma A.1 shows [^5] While Lemma A.1 is only stated for complex coefficient, we can embed our coefficient ring $A\llbracket y\rrbracket[y^{-1}]$ into $\mathbb{C}$ before using the lemma. that $g(w^{-1})$ is the compositional inverse of $f(y)$. That is, we have $f(g(w^{-1}))=w$, or equivalently $f(g(w))=w^{-1}$.
We write $g(w)=\sum_{k>0}g_{k}w^{k}$ and have $g_{k}=[w^{k}]g=\frac{1}{k}[y^{-1}]f^{k}$. By (4.1), $g_{k}\in A_{k-1}$. This allows us to specialize $w$ to any fixed value $c$ and the resulting series $g(c)$ gives a well-defined element in the formal completion $\widehat{A}$.
We set $w=-1$. Then $f(g(w))=w^{-1}$ becomes $f(g(-1))=-1$. It is the substitution of an element of $\widehat{A}$ for $y$ in an element of $A\llbracket y\rrbracket[y^{-1}]$ and reads
(4.2) $f\left(-\exp\left({\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]f^{k}}\right)\right)=-1.$
We recognize that $\widetilde{y}=\frac{1}{g(-1)}$ and the assertion follows from (4.2) by inserting the definition of $f$ that gives $\theta(y)=f(y^{-1})$.
By Lemma A.1, the compositional inverse $g(w)$ uniquely determines $\theta(y)$. We have argued above that $g_{k}\in A_{k-1}$. Hence, the $w$-exponent in $g$ only captures the degree of $g_{k}$ in the graded ring $A$. In turn, we can recover $g(w)$ from any fixed value of $w$. In particular, we can recover $g(w)$ from $\tilde{y}=\frac{1}{g(-1)}$ and this determines $\theta(y)$ under the requirement that $\theta(y)$ is of the given form. ∎
Using the quantum A-period $a(z,q)$ from Definition 2.14, in Definition 2.26 we defined a natural $q$-refinement of the open and closed mirror map by
@@
The ring $A:=\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$ is graded by $\deg(z^{\beta})=\beta\cdot E$. By Proposition 3.8, $\vartheta_{1}(y,z,q)_{\infty}$ is a formal Laurent series over $A$ of the form $\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$. Then Proposition 4.1 shows that the insertion of
@@
$M(Q,q)=1+\sum_{\beta}\sum_{g\geq 0}\frac{1}{\beta\cdot E-1}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$

### v2

@@ 2.2. Definition of the mirror curve
We view the mirror curve as being fibered over Spec [$\text{Spec }\mathbb{C}[\textup{NE}(\mathbb{P})]$ via the closed moduli $z^{\beta}$, and we write $p(x,y,z)$. Write $p(x,y,z)=1-f(x,y,z)$. We will give an explicit definition of the mirror curve by specifying the potential $f$.
@@
$\sum_{i=1}^{r+2}Q_{ik}m_{k}=0,\quad k=1,\ldots,r,$
can be packed into the charge matrix $Q=(Q_{ik})_{i,k}$. Using Gauss elimination we can transform the first $r$ columns of $Q$ into the identity matrix, so that $Q_{ik}=\delta_{ik}$ for $i,k\leq r$. Since $\mathbb{P}_{\Sigma}$ is smooth, the ray generators of $\rho_{r+1}$ and $\rho_{r+2}$ span the lattice, so all entries of $Q$ are integers. The rows of $Q$ define generators $\beta_{1},\ldots,\beta_{r}$ of the group $H_{2}(\mathbb{P}_{\Sigma},\mathbb{Z})$ by intersections with toric divisors
@@
Here we consider the indices modulo $r+2$. For an integer $l\geq 0$ and ordered tuples of integers $0\leq i_{1}<\ldots<i_{l}\leq m_{k}$ and $0\leq j_{1}<\ldots<j_{l}\leq n_{k}$ define
$\beta_{k,(i_{1},\ldots,i_{l}),(j_{1},\ldots,j_{l})}=\beta_{k}+\sum_{p=1}^{l}\sum_{q=0}^{i_{p}-1}\tilde{\beta}_{k-q}+\sum_{p=1}^{l}\sum_{q=1}^{j_{p}-1}\tilde{\beta}_{k+q}.$
@@
Define the (toric) potential of $\mathbb{P}_{\Sigma}$, resp. $\mathbb{P}$, by
@@
    The relation obtained from specializing the closed moduli $z^{\beta}$ by pairing with the polarization was already known by [CPS], Theorem 8.2, which amounts to saying that $\binom{\ell}{m_{k}}=\binom{\ell}{n_{k}}$ counts the summands in the $x^{a_{k}}y^{b_{k}}$-coefficient of $\vartheta_{1}$. Lemma 3.6 extends the agreement in part (1) of this remark by giving the correct closed moduli.
@@
Consider $\mathbb{P}^{1}\times\mathbb{P}^{1}$. A fan $\Sigma$ for $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is generated by $(-1,0)$, $(0,-1)$, $(1,0)$ and $(0,1)$. Again, all ray generators are vertices of $\Delta^{\star}$. The construction above gives
@@
Consider the Hirzebruch surface $\mathbb{P}_{\Sigma}=\mathbb{F}_{2}$. It admits a $\mathbb{Q}$-Gorenstein deformation to $\mathbb{P}^{1}\times\mathbb{P}^{1}$. The fan $\Sigma$ is generated by $(-1,0)$, $(-2,-1)$, $(1,0)$ and $(0,1)$. The construction above gives
@@ 2.4. The quantum A-period
According to [CKYZ], §6, one (classical) period to consider is the integral of $\log(p)$ over the torus $|x|=|y|=1.$ We call this the A-period. Note that with $p=-f(x,y,z)$ we have $\log(p)=-\sum_{n\geq 1}\tfrac{1}{n}f^{n}$, and each term $f^{n}$ is a Laurent polynomial in $x$ and $y$. Recall from Section 1.4 that for such Laurent series we write ${\sf Const}(F)$ for the constant term — then the integral picks up the constant term ${\sf Const}(\log(p))=-\sum_{n\geq 1}\tfrac{1}{n}{{\sf Const}}(f^{n})$. We use this fact to generalize the definition of period to the non-commutative case, with non-commuting variables $X$ and $Y$ obeying $YX=qXY.$
@@ 2.5. Wavefunctions and the A-period
It is shown in [ACDKV], Section 2.2 that the one-form $\log Yd\log X$ acquires a pole at the point of brane insertion with residue $\hbar$ (recall $\hbar=g_{s}=\lambda$ is the string coupling) – equivalently $W$ changes by $\hbar$ times a period.333 Quantum mirror curves play a central role in the Topological Strings/Spectral Theory duality of [GHM], where the spectrum of the operator defining the mirror curve is studied. Further period computations can be found in many other works — see, e.g., [DH, HMMO, HK, HW, DL], for example.
@@
See Section 2.6 for the computation of $Y(X)$ for $(\mathbb{P}^{2},E)$. We make no claims about the existence of other solutions.
@@
Let $Y(X)$ be the solution to $p(X,Y)=0$ from Proposition 2.19, and let $\Psi(X)$ be the wavefunction defined by $p\cdot\Psi=0$. Then
@@ 2.6. Computing the quantum A-Period
Remark 2.23.
Tropical methods have also been made to work recently by Singh and Zeitlin for different quantum integrable systems, [SZ].
@@ 2.7. Equality of definitions for the quantum A-period
Recall from Section 2.2 that $p(X,Y)=1-f(X,Y)$ gets quantized to an operator which acts on a function $\Phi(X)$ by $X\cdot\Phi(X)=X\Phi(X)$ and $Y\cdot\Phi(X)=\Phi(qX)$, and the wavefunction $\Psi(X)$ is defined by $p\cdot\Psi=0$. We have defined
@@
Substituting $X=xt$ and $Y=yt$, we get $f(X,Y,z)=tf(x,y,z)|_{z^{\beta}=t^{\beta\cdot E}z^{\beta}}$. E.g., for $\mathbb{P}^{2}$ we have $f=X+Y+zX^{-1}Y^{-1}=t(x+y+t^{3}zx^{-1}y^{-1})$. Applying $D$ as in Section 1.4, the claim is equivalent to
@@
Note that $g(t,y):=(1-tf)^{-1}\in(R[y]\llbracket y^{-1}\rrbracket)\llbracket t\rrbracket$ for $R=\mathbb{C}[\textup{NE}(X)][x]\llbracket x^{-1}\rrbracket$. By Convention 2.5, we can assume that $f(X,Y)$ does not contain any terms of $Y$-order $>1$. In other words, $\operatorname{ord}_{Y}(f)=1$, with Definition B.2. By Lemma B.3, $y(x,f)$ is a Laurent series of order $1$ in $f$. Then we can use Lemma B.6 to get
@@ 2.8. Natural $q$-refinements of mirror maps
where the last equality holds for every non-zero class $\beta\in\operatorname{NE}(\mathbb{P})$. We call the series $M(Q,q)$ a $q$-refined open mirror map. We will give it an enumerative interpretation in Theorem 4.2.
@@ 3. Quantum periods, broken lines and descendant invariants
In this section, we give a combinatorial construction of the quantum period (Definition 2.14) via broken lines in a consistent scattering diagram. In Section 4 we will use this description to prove that the primitive theta function, defined using broken lines, equals the open mirror map, also after $q$-refinement. Moreover, we discuss a relationship with descendant Gromov-Witten invariants with $\lambda$-class insertions.
@@ 3.1. Toric degenerations and scattering diagrams
Let $\mathbb{P}$ be a smooth projective Fano surface and let ${E}$ be a smooth anticanonical divisor, so that $(\mathbb{P},{E})$ is a smooth log Calabi-Yau pair. As in Section 2.1, if $\mathbb{P}$ is not toric, consider a $\mathbb{Q}$-Gorenstein degeneration of $\mathbb{P}$ to a Gorenstein toric Fano variety $\mathbb{P}_{\Sigma^{\prime}}$ or a smooth toric semi-Fano variety $\mathbb{P}_{\Sigma}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the ray generators of $\Sigma^{\prime}$ (or $\Sigma$) and let $\Delta$ be the polar dual polytope. Then $\Delta$ is the toric polytope of $\mathbb{P}_{\Sigma^{\prime}}$. The lattice points of $\Delta$ define an embedding of the projective toric variety $\mathbb{P}_{\Sigma^{\prime}}$ into $\mathbb{P}^{|\Delta\cap\mathbb{Z}^{2}|-1}$. The polytope $\Delta$ is reflexive. In particular it contains a unique interior lattice point.
@@ 3.2. Theta functions and quantum periods
We recall the definitions of broken lines and theta functions from [CPS][GHS], to which we refer for further details.
@@
Recall from Section 1.4 that we defined an operation $D$ on $\mathbb{C}[\textup{NE}(\mathbb{P})]$ by $Dz^{\beta}=(\beta\cdot D)z^{\beta}$.
@@
where $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ is the tropical analogue of $R_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$, defined as a count (with $q$-refined multiplicity) of rational tropical curves on the dual intersection complex $B$ of $(\mathbb{P},E)$ with an unbounded leg of weight $\beta\cdot E-1$ through a prescribed point, an unbounded leg of weight $1$ and no other unbounded legs. By [Gra2], Theorem 6.2, we have
@@
For $(\mathbb{P}^{2},E)$, Figure 3.3 shows the broken lines giving the first few terms of
@@ 3.3. Quantum descendant invariants
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the top Chern classes of the Hodge bundle.
@@ 4. The $q$-refined open mirror map is the quantum theta function
We prove the identification of the quantum theta function asymptotic broken line expansion and the $q$-refined open mirror map in this section. As a consequence, we obtain an enumerative interpretation of the $q$-refined open mirror map.
@@
Let $A=\bigoplus_{k\geq 0}A_{k}$ be a finitely generated graded algebra over $A_{0}$ and let $\widehat{A}$ denote the formal completion of $A$ in the ideal $\bigoplus_{k\geq 1}A_{k}$. Let $\theta(y)\in A\llbracket y^{-1}\rrbracket[y]$ be a formal Laurent series of the form $\theta(y)=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$.
  1. (1)
    The series $\eta(w)=w\exp\left({-\sum_{k>0}\frac{1}{k}[y^{0}]\theta^{k}w^{-k}}\right)$ gives a well-defined element of $A\llbracket w^{-1}\rrbracket[w]$ of the form $\eta(w)=w+\sum_{k\geq 0}b_{k}w^{-k}$ that is a compositional inverse of $\theta(y)$, i.e., it satisfies $\theta(\eta(w))=w.$
  2. (2)
    The series $\widetilde{y}=-\exp\left({-\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]\theta^{k}}\right)$ gives a well-defined element of $\widehat{A}$ that satisfies $\theta(\tilde{y})=-1.$ Moreover, $\tilde{y}$ uniquely determines the coefficients $a_{k}$, hence it determines a function $\theta(y)$ of the given form.
@@
Define $f(y)=\theta(y^{-1})$ and let $A_{r}$ be the degree $r$ graded piece of $A$. Then $f(y)$ is a formal Laurent series in $y$ of the form $f(y)=y^{-1}+\sum_{k\geq 0}f_{k}y^{k}$ such that $f_{k}\in A_{k+1}$. Recall from Section 1.4 that we write $[y^{k}]f$ for the coefficient of $y^{k}$ in $f$. The power series $h(y):=yf(y)=1+\sum_{k>0}f_{k-1}y^{k}\in A\llbracket y\rrbracket$ has the convenient property $[y^{k}]h^{l}\in A_{k}$ for every $l\geq 0$. In particular, $[y^{k}](yf)^{k}\in A_{k}$, so $[y^{0}]f^{k}\in A_{k}$ and
@@
We apply Proposition B.12 to $h$ to have
@@
Comparing the very last expression with Lemma B.1 shows [^6] While Lemma B.1 is only stated for complex coefficient, we can embed our coefficient ring $A\llbracket y\rrbracket[y^{-1}]$ into $\mathbb{C}$ before using the lemma. that $g(w^{-1})$ is the compositional inverse of $f(y)$. That is, we have $f(g(w^{-1}))=w$, or equivalently $\theta(g(w^{-1})^{-1})=w$. From the definitions of $g(w)$ and $\eta(w)$ we see that $\eta(w)=g(w^{-1})^{-1}$, proving (1).
We write $g(w)=\sum_{k>0}g_{k}w^{k}$ and have $g_{k}=[w^{k}]g=\frac{1}{k}[y^{-1}]f^{k}$. By (4.1), $g_{k}\in A_{k-1}$. This allows us to specialize $w$ to any fixed value $c$ and the resulting series $g(c)$ gives a well-defined element in the formal completion $\widehat{A}$. We set $w=-1$. Then $\theta(g(w^{-1})^{-1})=w$ becomes $\theta(g(-1)^{-1})=-1$. It is the substitution of an element of $\widehat{A}$ for $y$ in an element of $A\llbracket y^{-1}\rrbracket[y]$. We recognize that $\widetilde{y}=g(-1)^{-1}$, giving $\theta(\widetilde{y})=-1$.
By Lemma B.1, the compositional inverse $g(w)$ uniquely determines $\theta(y)$. We have argued above that $g_{k}\in A_{k-1}$. Hence, the $w$-exponent in $g$ only captures the degree of $g_{k}$ in the graded ring $A$. In turn, we can recover $g(w)$ from any fixed value of $w$. In particular, we can recover $g(w)$ from $\tilde{y}=\frac{1}{g(-1)}$ and this determines $\theta(y)$ under the requirement that $\theta(y)$ is of the given form. This proves (2), completing the proof. ∎
Using the quantum A-period $a(z,q)$ from Definition 2.14, in Definition 2.27 we defined a natural $q$-refinement of the open mirror map by
@@
The ring $A:=\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$ is graded by $\deg(z^{\beta})=\beta\cdot E$. By Proposition 3.8, $\vartheta_{1}(y,z,q)_{\infty}$ is a formal Laurent series over $A$ of the form $\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$. Then Proposition 4.1, (2) shows that the insertion of
@@
$M(Q,q)=1+\sum_{\beta}\sum_{g\geq 0}\frac{(-1)^{\beta\cdot E}}{\beta\cdot E-1}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$